The common-ion effect is an application of Le Châtelier's principle, which states that an equilibrium system that is stressed will work to alleviate that stress and reestablish equilibrium.

What is the pH change of a 0.250 M solution of citric acid (pKa=4.77) if citrate is added to a concentration of 0.135 M with no change in volume?

Express the difference in pH numerically to two decimal places.
Answer #1

pH = 1/2 [pKa - log C]

    = 1/2 [4.77 - log 0.250]

    = 2.69

new pH

pH = pKa + log [citrate / citric acid ]

       = 4.77 + log (0.135 / 0.250)

       = 4.50

pH difference = 4.50 -2.69

                       = 1.81

difference in pH = 1.81